OpenStudy (skarlettrose186):

Using the formula W = mg, how many milliliters of water with a density of 1g/mL are required to weigh 0.75 newtons and g = 9.81 m/s2? Round to the nearest tenth. (Note: The mass will be in kg in your answer, thus convert to g and then to mL.)

OpenStudy (skarlettrose186):

@Awolflover1 @IrishBoy123

OpenStudy (skarlettrose186):

@ILovePuppiesLol

OpenStudy (zach298):

w = mg 0.75N = m * 9.81 m = 0.75 / 9.81 = 0.0765 kg The mass in grams is 0.0765 * 1000 = 76.5g At a density of 1 g/ml, 76.5g ÷ 1g/ml = 76.5ml